# Covariances for Free: Exploiting Mean Distributions for Federated Learning with Pre-Trained Models

### v1

@@ Appendix H Additional Experiments
Impact of using pre-trained models. To quantify impact of using pre-trained models we performed experiments using a randomly initialized model and show in Table 5 that federated training using a pre-trained model significantly outperforms a randomly initialized model using standard methods like FedAvg and FedAdam on CIFAR-10 and CIFAR-100.
Table 5: Impact of using pre-trained SqueezeNet network with different federated learning methods on CIFAR100. We show the total communication cost (in MB) from all clients to server. We train 100 clients with 30 clients per round for 200 rounds in non-iid settings with dirichlet distribution of 0.1. When starting from random initialization (no pre-training), we train for 400 rounds.
CIFAR10 CIFAR100
FedAvg $\times$ 37.3 74840 23.9 79248
FedAdam $\times$ 60.5 74840 44.3 79248
FedAvg $\checkmark$ 84.7 37420 56.7 39624
FedAdam $\checkmark$ 85.5 37420 62.5 39624
Appendix I Convergence Analysis
In our work, we claim that FedCOF initialization achieves faster and better convergence based on our empirical results (Fig. 7) using multiple datasets. We propose how to initialize the classifiers before performing federated optimization methods like FedAvg (McMahan et al., 2017) and FedAdam (Reddi et al., 2020) which have already established the theoretical guarantees of convergence in their respective works. Unlike gradient-based FL methods, our method is training-free. Similar to Fed3R (Fanì et al., 2024) and FedNCM (Legate et al., 2023a), the proposed FedCOF does not depend on assumptions like bounded variance of stochastic gradients or smoothness of clients objectives. While we do not propose any federated optimization step, we propose a training-free method that can be also used for initializing federated fine-tuning. We would also like to highlight that all existing works in Federated learning with pre-trained models (Tan et al., 2022b; Nguyen et al., 2020; Fanì et al., 2024; Legate et al., 2023a; Chen et al., 2022; Qu et al., 2022; Shysheya et al., 2022) focus only on empirical observations assuming that the theoretical guarantees of existing federated optimization methods holds true when using pre-trained models. A more exhaustive study on convergence analysis for FL with pre-trained models would be an interesting direction to explore in future works.
Appendix J Bias of the Estimator with non-iid Client Features
In Appendix C we showed that, under the assumption that the per-class features are iid across clients, the proposed estimator is an unbiased estimator. In this section, we theoretically quantify the bias when the i.i.d assumption is violated.
Under the i.i.d. assumption, the single class features assigned to clients can be treated as random samples from the same population distribution with mean $\mu_{c}$ and covariance $\Sigma_{c}$. For simplicity, focusing on a single class and dropping the class subscript $c$, the population distribution has mean $\mu$ and covariance $\Sigma$. As a result, recalling equation 16, we can write:
To compute the expectation of the estimator $\mathbb{E}[\hat{\Sigma}]$, we follow the same procedure used to prove proposition in Appendix C up to Eq. 15:
Substituting Equations 24 and 22 into Equation 23, and recalling that $N=\sum_{k=1}^{K}n_{k}$, we obtain:
where in the last step we used that $\sum_{k=1}^{K}n_{k}\mu_{k}=N\mu$.
The bias of the estimator is thus given by:
Note that if each client population covariance $\Sigma_{k}$ is equal to the global population covariance $\Sigma$, and the mean of each client $\mu_{k}$ is equal to the population mean, then the bias is zero (i.e., the estimator is unbiased). However, the bias formula reveals that when the distribution of a class within a client differs from the global distribution of the same class, our estimator introduces a systematic bias. This situation can arise in the feature-shift setting, in which each client is characterized by a different domain. In the next section, we evaluate FedCOF under the feature-shift setting to quantify how this bias affects performance in this specific scenario.
As a final note, we mention that we always assume the global distribution of a single class can be modeled with a distribution having a single mean and covariance (see Eq. 24). This is how our classifier operates. As future work, it could be beneficial to employ different types of classifiers that allow multiple class means and class covariances.
Appendix K Experiments on feature shift settings
Method Acc ($\uparrow$) Comm. ($\downarrow$)
FedNCM 65.8 0.3
Fed3R 81.9 39.6
FedCOF 74.1 0.3
FedCOF (2 class means per client) 76.5 0.6
FedCOF (10 class means per client) 78.8 3.1
Following (Li et al., 2021), we perform experiments with MobileNetv2 in a non-iid feature shift setting on the DomainNet (Peng et al., 2019) dataset. DomainNet contains data from six different domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch. We use the top 10 most common classes of DomainNet for our experiments following the setting proposed by (Li et al., 2021). We consider six clients where each client has i.i.d. data from one of the six domains. As a result, different clients have data from different feature distributions. We show in Table 6 how training-free methods perform in feature shift settings and the accuracy to communication trade-offs.
Appendix L Privacy Concerns on sharing Class-wise Statistics
Our method requires transmitting class-wise statistics to compute the unbiased estimator of the population covariance (Eq. 14) and classifier initialization, similar to other methods in federated learning (Legate et al., 2023a; Luo et al., 2021). In general, transmitting the class-wise statistics may raise privacy concerns, since each client could potentially expose its class distribution. Inspired by differential privacy (Dwork et al., 2006), we propose perturbing the class-wise statistics of each client with different types and intensities of noise, before transmission to the global server. This analysis allows us to evaluate how robust FedCOF is to variations in class-wise statistics and whether noise perturbation mechanisms can effectively hide the true client class statistics. Specifically, we propose perturbing the class-wise statistics as follows:
$\widetilde{n}_{k,c}=\max(n_{k,c}+\sigma^{\text{noise}}_{\epsilon},0)$ (26)
We consider three types of noise:
In Figure 16, we show that the performance of FedCOF is robust with respect to the considered noise perturbation, varying the intensity of $\epsilon\in\{0.1,0.3,0.5,0.7,0.9\}$. These results suggest that a differential privacy mechanism can be implemented to mitigate privacy concerns arising from the exposure of client class-wise frequencies. In Figure 17, we provide a qualitative overview of how the proposed Laplacian and uniform noise perturbation affect class-wise distributions.
Appendix M Dataset Details
Apart from CIFAR100, we perform experiments with ImageNet-R (Hendrycks et al., 2021) which is an out-of-distribution dataset and proposed to evaluate out-of-distribution generalization using ImageNet pre-trained weights. It contains data with multiple styles like cartoon, graffiti and origami which is not seen during pre-training. We also consider fine-grained datasets like CARS and CUB200 for our experiments. Finally, we also use iNaturalist-Users-120k (Hsu et al., 2019) dataset in our experiments, which is a real-world, large-scale dataset (Van Horn et al., 2018) proposed by (Hsu et al., 2019) for federated learning and contains 120k training images of natural species taken by citizen scientists around the world, belonging to 1203 classes spread across 9275 clients. In datasets like ImageNet-R and CARS, we also face class-imbalanced situations where there is a significant class-imbalance at the global level.

### v3

@@ I.2 FedCOF with Secure Aggregation Protocols
Figure 9: (left) Analysis of FedCOF performance with multiple class means per client on ImageNet-R. (center) Total number of means per class on average that are used to estimate the covariance for FedCOF in Figure˜9 (left). (right) Performance comparison of FedCOF with full, diagonal, and spherical covariance matrix communication.
Appendix L Additional Ablations
Sampling multiple class means. We perform multiple class means sampling per client using ImageNet-R and show in Figure˜9 (left) that using FedCOF with more class means shared from each client improves the performance. We also show in Figure˜9 (center) the total number of means used per class on an average in Figure˜9 (left) to perform the covariance estimation. The number of means used to estimate each class covariance is less than the total number of clients due to the class-imbalanced or dirichlet distribution used to sample data for clients. This is due to the fact that not all classes are present in all clients.
In most settings we observe that the largest performance improvement occurs when increasing the number of class means per client in scenarios with fewer clients. For instance, with only 10 clients, accuracy improves from 33.1% to 37.0% when increasing from 1 to 4 means per client. In contrast, when 50 clients are available, the improvement is marginal (from 37.4% to 37.8%). This trend is consistent with the theoretical insights in Appendix˜E: when fewer clients are available, sampling more means per client increases the number of independent statistics, thereby reducing variance and improving the quality of the covariance estimate.
Communicating diagonal or spherical covariances. While communicating diagonal or spherical covariances (mean of the diagonal covariance) from clients to server and then estimating the global class covariance from them can significantly reduce the communication cost, such estimates of global class covariance is poor compared to FedCOF. We show in Figure˜9 (right) that FedCOF outperforms these covariance sharing baselines when communicating spherical or diagonal covariances.
Table 14: Ablation showing the impact of using shrinkage in FedCOF using a pre-trained SqueezeNet.
Dataset $\gamma=0$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=1$ $\gamma=10$
ImageNet-R 36.53 36.98 36.96 37.25 36.07
CUB200 51.08 51.07 51.81 53.57 53.50
Impact of Shrinkage. We analyze the impact of shrinkage on the estimated class covariances in FedCOF when using a pre-trained SqueezeNet in Table˜14. We use a shrinkage $\gamma=1$ for our experiments with SqueezeNet and ViT-B/16. We observe that shrinkage yields marginal improvement for ImageNet-R and a bit more significant improvement in accuracy of 2.5% on CUB200. This can be attributed to the few-shot settings where the covariance estimation is not very good due to the lack of data and thus fewer clients having access to each of the classes. The use of shrinkage in FedCOF stabilizes and improves the covariance estimation leading to improved accuracy especially in few-shot settings.
Severe imbalance settings. We further evaluate FedCOF in settings with more severe class imbalance or heterogeneity across clients (Dirichlet distribution with $\alpha=0.05,0.01$) in Table˜15 and show that the performance of FedCOF drops a bit with very severe heteregeneity as expected.
Table 15: Analysis of FedCOF performance across settings with varying heterogeneity.
Class means shared ImageNet-R (100 clients) Cars (100 clients)
per client $\alpha=0.5$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.5$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
1 38.4 37.3 36.4 33.5 45.0 44.5 43.1 39.9
2 38.2 37.8 37.4 35.9 44.9 44.5 43.8 42.5
3 38.1 37.5 37.5 36.7 44.9 44.7 44.2 43.2
4 38.3 37.7 38.2 37.4 45.1 44.9 44.4 43.6
Table 16: Total means shared per class on average (the number of clients each class is present in, on average).
Dataset $\alpha=0.5$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
ImageNet-R 36.1 17.4 11.7 4.0
Cars 24.5 13.4 9.6 3.6
We show the impact of class imbalance in different settings in Table˜16. Using the most extreme setting ($\alpha=0.01$), each class is present on 4 clients on an average. Although the extreme settings are less unrealistic, we analyze and evaluate FedCOF in those settings. For instance, the drop in performance of FedCOF on Cars from 44.5 to 39.9 is due to the fact that the global covariance is estimated from around 3.6 sample means instead of around 13.4 sample means. This scenario faces the same issue as fewer clients due to the severe class imbalance. Here, we show that sampling multiple class means per client improves the performance in the extreme settings mitigating the bias in these settings.